OpenStudy (anonymous):

Write the series in summation notation: -3 + 6 - 9 + 12 - 15

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} 3n (-1)^{n}\]

OpenStudy (anonymous):

what about 1+1/4+1/9+1/16+1/25

OpenStudy (anonymous):

that series in summation notation

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ 1 }{ n ^{2} }\]

OpenStudy (anonymous):

thx:)

OpenStudy (anonymous):

pleasure